A 60 W fluorescent bulb has resistance of about 12 Ω when cold (30˚C) 

and 120 Ω when on (hot). Estimate the temperature of the filament when 

hot assuming an average temperature coefficient of resistivity. Where α = 0.0080(C˚)-1




1
Expert's answer
2021-03-26T11:38:25-0400
"R=R_0(1+\alpha (T-T_0))\\120=12(1+0.008 (T-30))\\T=1155\degree C"